An elephant (weight 50000 N) sits on the large pla
Answer #1

a)

F1 = Force on the large plate = 50000 N

A1 = area of large plate = 10 m2

F2 = Force on the small plate

A2 = area of small plate = 0.01 m2

Using pascal's law

F1 / A1 = F2 / A2

50000/10 = F2 / 0.01

F2 = 50 N

b)

F1 d1 = F2 d2

50000 x d1 = 50 x 1

d1 = 0.001 m